OpenStudy (anonymous):

How do you find the slop of 10x-5y=30

OpenStudy (katrinakaif):

Lets change this to y = mx + b format

OpenStudy (katrinakaif):

To do so, it becomes : 10x - 5y = 30 -5y = -10x + 30 y = 2 x - 6

OpenStudy (katrinakaif):

Therefore, 2 is your slope.

OpenStudy (anonymous):

I dont understand, why just 2 is the slope then?

OpenStudy (anonymous):

Would that mean 6=y?

OpenStudy (katrinakaif):

The equation is in y = mx +b form y = 2x - 6 m is your slope therefore, 2

OpenStudy (katrinakaif):

You can't solve these equations since you have 2 variables, x and y

OpenStudy (anonymous):

Katty is correct. I only told the answer to her.

OpenStudy (anonymous):

10x-5y=30 to find the slope First we need to solve for 'y' that given equation subtract 10x on both sides -5y = -10x +30 divide by -5 on both sides y = 2x - 6 We know the slope intercept form y = mx + b where m = slope b = y-intercept y = 2x - 6 compare this equation to slope intercept form y = mx + b Therefore , slope = m = 2 , y-intercept = b = -6 Slope = 2 y-intercept = -6